40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 03:08:57 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec DRFPMI_locked.pdf
    Reply  Thu Oct 15 04:14:14 2015, ericq, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Thu Oct 15 10:59:12 2015, Koji, Update, LSC, DRFPMI Locked for 20 sec 
    Reply  Sat Oct 17 18:55:07 2015, rana, Update, LSC, DRFPMI Locked for 20 sec 
       Reply  Mon Oct 19 15:23:22 2015, ericq, Update, LSC, Longer DRFPMI lock second_DRFPMI_15Oct2015.png
          Reply  Tue Oct 20 11:24:29 2015, ericq, HowTo, LSC, How to DRFPMI 
Message ID: 11701     Entry time: Tue Oct 20 11:24:29 2015     In reply to: 11698
Author: ericq 
Type: HowTo 
Category: LSC 
Subject: How to DRFPMI 

Herein, I will describe the current settings and procedures used to achieve the DRFPMI lock, cobbled together from scripts, burts and such. 


Initial Alignment

  1. With arms POX/POY locked, run dither alignment servos. Set transmon QPD offsets here
  2. Restore "PRMI Carrier" configuration, run BS and PRM dither alignment servos simultaneously. (Note: this sacrifices some X arm alignment for better dark port alignment. In practice no appreciable loss of TRX is observed)
  3. Misalign PRM, align SRM and tune SRM alignment by eye while looking at AS camera. 
  4. Restore POX/POY arm lock, lock green to arms, check that powers are high enough and align if neccesary.

Initial Configuration

CARM, DARM

For CARM and DARM, the A channels are used for the ALS signals, whereas the B channels are used for blending the RF signals. 

ALS

  • BEATX and BEATY, I and Q channels: +0dB Whitening Gain, Whitening Filters ON
  • Green beatnotes somewhere between 20-80MHz, following sign convention of temperature slider UP makes beat freq go UP.  Check spectrum of PHASE_OUT_HZ vs references in ALS_outOfLoop_Ref.xml. The locking script automatically sets the correct phase tracker gain, so no need to adjust manually.
  • CARM_A = -1.0 x ALSX + 1.0 x ALSY, G=1.0
  • DARM_A = 1.0 x ALSX + 1.0 x ALSY, G=1.0

RF

  • CM Board: REFL11 I daugher board output -> IN1, IN1 Enabled, -32dB input gain, 0.0V offset, all boosts off, AO polarity positive, AO gain +0dB
  • MC Board: IN2 disabled, -32dB input gain
  • CM_SLOW: +0dB Whitening Gain, Whitening ON, LSC-CM_SLOW_GAIN = -5e-4 (Though, it would be good to reallocate this gain to the input matrix element)
  • CARM_B = 1.0 x CM_SLOW, FM4 FM10 ON, G=0 (FM4 = LP700 for AO crossover stability, FM10 = 120:5k for coupled cavity pole compensation)
  • AS55: +9dB Whitening Gain, Whitening filters manual, Demod angle -37.0
  • DARM_B = -1e-4 x AS55 Q, G=0

DRMI 3F

For the DRMI, the A channels are used for the 1F signals, whereas the B channels are used for the 3F signals. The settings for transitioning to 1F after locking the DRFPMI have not yet been determined. 

These settings are currently saved in the DRMI configurator, but the demod angles are set for DRFPMI lock, so the settings don't reliably work for misaligned arms. 

  • REFL33: +30dB Whitening Gain, Whitening filters trigger on DRMI lock, Demod angle: 136.0
  • REFL165: +24dB Whitening Gain, Whitening filters trigger on DRMI lock, Demod angle: -111.0
  • POP22: +15dB Whitening Gain, Whitening filters OFF, Demod angle: -114.0
  • AS110: +36dB Whitening Gain, Whitening filters OFF, Demod angle: -116.0
  • POPDC: +0dB Whitening Gain, Whitening filters OFF (used as a supplemental trigger signal when CARM and DARM are buzzing and POP22 fluctuates wildly)
  • MICH_B = 6.0 x REFL165Q, offset = 15.0
  • PRCL_B = 5.0 x REFL33I, offset = 45.0
  • SRCL_B = -0.6 x REFL165I + 0.24 x REFL33 I, offset=0

The REFL33 element in SRCL_B is to reduce the PRCL coupling, was found empirically by tuning the relative gains with the arms misaligned and looking at excitation line heights. The offsets were found by locking the DRMI on 1F signals with arms misaligned, and taking the average value of these 3F error signals.  

Servo filter configuration

The CARM and DARM ALS settings are largely scripted by scripts/ALS/Transition_IR_ALS.py, which takes you from arms POX/POY locked to CARM and DARM ALS locked. The DRMI settings are usually restored from the IFO_CONFIGURE screen. 

  • CARM: FM[1, 2, 3, 5, 6] , G=4.5, Trigger forced on, no FM triggers, output limit 8k
  • DARM: FM[1, 2, 3, 5, 6] , G=4.5, Trigger forced on, no FM triggers, output limit 8k
  • MICH: FM[4, 5], G= -0.03, Trigger POP22 I x 1.0 [50, 10], FM[2, 3, 7] triggered [50, 10], output limit 20k
  • PRCL: FM[4, 5], G= -0.003, Trigger POP22 I x 1.0 [50, 10], FM[1, 2, 8, 9] triggered [50, 10], output limit 8k
  • SRCL: FM[4, 5], G= -0.4, Trigger AS110 Q x 1.0 [500, 100], FM[2, 7, 9] triggered [500, 100], output limit 15k

Actuation Output matrix

  • MC2 = -1.0 x CARM
  • ETMX = -1.0 x DARM
  • ETMY = 1.0 x DARM
  • BS = 0.5 x MICH
  • PRM = 1.0 x PRCL - 0.2655 MICH
  • SRM = 1.0 x SRCL + 0.25 MICH (The mich compensation is very roughly estimated)

Locking Procedure

When arms are POX/POY locked, and the green beatnotes are appropriately configured, calling scripts/DRFPMI/carm_cm_up.sh initiates the following sequence of events:

  • Turn ON MC length feedforward and PRC angle feedforward
  • Set ALS phase tracker UGFs by looking at I and Q magnitudes
  • Set LSC-ALSX and LSC-ALSY offsets by averaging, ramp CARM+DARM gains up, XARM+YARM gains down, engage CARM+DARM boosts, now ALS locked
  • Move CARM away from resonance, offset = -4.0 (DRMI locks quicker on this side for whatever reason)
  • Restore PRM, SRM alignment. Set DRMI A FM gains to 0, B FM gains to 1.0. Enable LSC outputs for BS, PRM, SRM
  • When DRMI has locked, add POPDC trigger elements to DRMI signals and transition SRCL triggering to POP22I. NB: In the c1lsc model, the POPDC signal incident on the trigger matrix has an abs() operator applied to it first. 
    • MICH Trig = 1.0 x POP22 I + 0.5 x POPDC, [50, 10]
    • PRCL Trig = 1.0 x POP22 I + 0.5 x POPDC, [50, 10]
    • SRCL Trig = 10.0 x POP22 I + 5 x POPDC, [500, 100]
  • Reduce POX, POY whitening gains from their nominal +45dB to +0dB, so there aren't railing channels making noise in the whitening chassis and ADCs
  • DC couple ITM oplevs (average spot position, set FM offset, turn on DC boost filter, let settle)
  • With an 8 second ramp, reduce CARM offset to 0 counts. 
  • MANUALLY adjust CARM_A and DARM_A offsets to where CARM_B_IN and DARM_B_IN are seen to fluctuate symetrically around their zero crossing.
    • Note: Last week, this adjustment tended to be roughly the same from lock to lock, unlike the PRFPMI which generally didn't need much adjustment. Also, by jumping from CARM offset of -0.4 to 0.4, it could be seen that the zero crossing in  CARM_B aka CM_SLOW aka REFL11 had some offset, so CARM_B_OFFSET was set to 0.005, but this may change. 

When CARM and DARM are buzzing around true zero, powers maximized:

  • CARM and DARM FM1 (18,18:1,1 boosts) OFF
  • CARM_B_GAIN 0.0 -> 1.0, FM7 ON (20:0 boost)
  • DARM_B_GAIN 0.0 -> 0.015, FM7 ON (20:0 boost) 
  • MC servo board IN2 ENABLE, IN2 gain -32dB -> -16dB
  • Turn ALL MC2 violin filters OFF (smoothen out AO crossover)
  • If stable, CM board IN1 gain -32dB -> -10dB (This is the overall CARM gain, the arm powers stabilize within the last few dB of this transition)
  • CARM_A_GAIN 1.0 -> 0.7
  • CARM_A FM9 ON (LP1k), sleep, FM 1 ON (1:20 deboost), sleep, FM 2 ON (1:20 deboost), HOLD OUPUT, CARM now RF only
  • DARM_B_GAIN 0.015 -> 0.02, sleep, DARM_A_GAIN 1.0 -> 0.0 (This may not be the ideal final DARM_B gain, UGF hasn't been checked yet)

IFO is now RF only!

  • Turn on transmon QPD servos.
  • Adjust comm/diff QPD servo offsets to correct any problems evident on AS/REFL cameras. This usually brings powers from ~100-120 to ~130-140. 

This is as far as we've taken the DRFPMI so far, but the CARM bandwidth is still only at a few kHz. Based on PRFPMI locking, the next steps will be:

  • CM BOARD +12dB or so additional IN1 gain, more AO gain may be needed to get crossover to final position of ~100Hz
  • MC2 viollin filters back on
  • CM boost(s) on
  • AS55 whitening on
  • Transition DRMI to 1F
ELOG V3.1.3-